An electron is moving around a hydrogen nucleus in a circular orbit the force between the two each

Grade:12th pass

2 Answers

Pooja
4 years ago
The force is given as ke/ r2
now mvr=nh/2π
mv2/r = ke2/r2
m(vr)2 = ke2r
(vr)2 = ke2r/m
 
 
(mvr)2 = n2h2/4π2
(vr)= n2h2/4π2m2 = ke2r/m
r = n2h2/4π2ke2m
F = ke2/r2
   = ke2/(n2h2/4π2ke2m)2
F = 16π4k3e6m/ n4h4
 
 
 
 
 
Khimraj
4 years ago
 
Dear student
Coulomb force = - ke² * (r vector)/r³
= - k* e² * (unit vector r)/r²
…...................................